OpenStudy (anonymous):

Solve the absolute value inequality: |5x + 3|>= 13 Answer x >= -3.2 or x <= 2 x <= -3.2 and x >= 2 x <= -3.2 or x >= 3.2 x <= -3.2 or x >= 2

OpenStudy (anonymous):

5x+3 <= -13 and 5x + 3 >=13 5x<=-16 and 5x>=10 x<=-16/5 and x>=2 x<=-3.2 and x>=2

OpenStudy (anonymous):

use four formulae (1) |x|<=a then -a<= x <= +a () |x|<a then -a < x < +a () |x|>=a then x<= -a and x>=+a () |x|> a then x< -a and x> +a in place of x you can substitute a bulky expression of x..

OpenStudy (riley):

Solve this one the same as the other problem. solve for x the normal way for one value. and then got back and solve for x switching the signs! ^_^ So: 5x+3 >= 13 5x >= 10 x>2 -5x-3 >= 13 -5x >=16 x<=-3.2 So your answers are x>=2 and x<=-3.2